OpenStudy (anonymous):

when 5 is added to 2 times a number the result is 45 Find the number

OpenStudy (phebe):

First you need to set up the problem

OpenStudy (phebe):

2*x+5=45 ^^^^^^ this is the right problem

OpenStudy (phebe):

All you need to do is solve it

OpenStudy (anonymous):

and what is the anwser

OpenStudy (nurali):

x - the number 5 + 2x = 45 2x = 45 - 5 2x = 40 x=?

OpenStudy (phebe):

If you solve 1x=40 You will get your answer

OpenStudy (anonymous):

thanks all of youll

OpenStudy (phebe):

just divide by 2 on both sides

OpenStudy (phebe):

Welcome!

OpenStudy (anonymous):

the answer is 20 but you probaly already got it